OpenStudy (anonymous):

Find the unit rate, and then use it to solve the problem. A trucking company makes 50 deliveries in 5 days. How long will it take for 400 deliveries? A. 8 days B. 10 days C. 20 days D. 40 days

OpenStudy (amistre64):

how many 50s are in 400?

OpenStudy (amistre64):

we have a ratio of 50:5 50k : 5k maintains that ratio 50k = 400, defines k for us ... and by extension 5k or just solve for 5k ... either way

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

cant you just divide it by 10?

OpenStudy (amistre64):

you can, hence the option to just solve it for 5k

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

arenti suspose to didve 10 by 400?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

the most basic approach is to determine the solution for 50k = 400 this defines k, which means we can know 5k

OpenStudy (anonymous):

50=400 5k=40

OpenStudy (anonymous):

i mean 4

OpenStudy (amistre64):

divide both sides by 50 50k/50 = 400/50 k = 40/5 multiply both sides by 5 5k = 40*5/5

OpenStudy (anonymous):

are you still lost?

OpenStudy (amistre64):

the account got terminated ... for reasons beyond my knowledge.